# Bayesian Quantum Amplitude Estimation

### v1

@@ Abstract
Quantum amplitude estimation is a fundamental routine that offers a quadratic speed-up over classical approaches. The original QAE protocol is based on phase estimation. The associated circuit depth and width, and the assumptions of fault tolerance, are unfavorable for near-term quantum technology. Subsequent approaches attempt to replace the original protocol with hybrid iterative quantum-classical strategies. In this work, we introduce BAE, a noise-aware Bayesian algorithm for QAE that combines quantum circuits with a statistical inference backbone. BAE can dynamically characterize device noise and adapt to it in real-time. Problem-specific insights and approximations are used to keep the problem tractable. We further propose an annealed variant of BAE, drawing on methods from statistical inference, to enhance statistical robustness. Our proposal is parallelizable in both quantum and classical components, offers tools for fast noise model assessment, and can leverage preexisting information. Additionally, it accommodates experimental limitations and preferred cost trade-offs. We show that BAE achieves Heisenberg-limited estimation and benchmark it against other approaches, demonstrating its competitive performance in both noisy and noiseless scenarios.
preprint: APS/123-QED
I Introduction
Quantum amplitude estimation (QAE) is a routine for estimating the measurement probability associated with a given substate for some wavefunction [1]. Notable applications are Monte Carlo integration [2] and estimation tasks [3], with applications in finance [4, 5, 6, 7, 8, 5], chemistry [9, 10, 11]) and machine learning [12, 13, 14, 15].
@@ I Introduction
On the other hand, the cost of a naive (deterministic) statistical approximation increases exponentially with the number of parameters. More efficient approximations are vital to ensure tractability; importantly, care must be taken lest they jeopardize scalability or correctness. In [29], a noise-aware approach to QAE based on Bayesian inference is proposed under a normality assumption. While this is a cost-effective approach for simple cases, it is not expected to generalize well to more complex ones, namely problems involving more elaborate noise models.
@@
In this work, we propose an approximate Bayesian algorithm that is capable of handling noise without making restrictive assumptions or incurring prohibitive costs. Problem tailored heuristics are used to cut down optimization costs while preserving the quantum advantage, bringing the optimization cost per iteration down to constant. By employing scalable and highly parallelizable statistical methods and other approximations, we further lower the classical processing costs without affecting the quantum advantage. The employed quantum circuits are of the simplest type presented in the literature; they do not require controlled or generalized versions of the Grover operator, nor any other qualitative complication as compared to Grover search.
@@
Our method offers quantifications of model merit with little extra cost, can incorporate available information such as past experimental data, and is flexible with respect to the cost trade-offs involved. For instance, one may trade in part of the quantum advantage for a lower classical cost, or for increased parallelism in the quantum sub-routine.
@@
To benchmark BAE, we numerically test its performance along with those of other QAE algorithms presented in the literature, and propose thorough and cost-effective methods to evaluate and analyze their performance.
@@
Additionally, inspired by a widely used method in statistics, we propose a variant of BAE where the quantum circuits are guided based on measures of statistical efficiency not directly related to the error or uncertainty. This further highlights the importance of this component, and offers a robust and systematic way to work with challenging models or data. We call this variant annealed Bayesian amplitude estimation, after the method it is inspired in, annealed likelihood estimation [30].
@@
The rest of the paper is organized as follows. Section II presents relevant background, including the original quantum algorithm, the classical counterpart, and alternative hybrid algorithms. In section III, BAE is presented. Section IV describes the numerical methods used to test the algorithms. The results of these tests are shown in section V. Finally, section VI discusses and summarizes the results, as well as directions for future research.
II Background
II.1 Amplitude estimation
@@
This type of wavefunction arises most commonly in the context of Grover search [31], where given a function $f(x):\{0,1\}^{n}\rightarrow\{0,1\}$, we want to find the pre-image of $1$ under $f$, i.e. $x$ s. t. $f(x)=1$. Additionally, we know how to prepare state (1), where $\left\lvert\psi_{1}\right\rangle$ is the subspace spanned by solution states and $\left\lvert\psi_{0}\right\rangle$ is its orthogonal complement in the Hilbert space. The probability of finding a solution upon measuring state (1) is then $a$. For that reason, $\left\lvert\psi_{1}\right\rangle$ is often called the good subspace, and $\left\lvert\psi_{0}\right\rangle$ is called the bad subspace.
@@
II.2 Quantum algorithm
Amplitude estimation is the task of learning $\theta$ or $a$. It can be shown that ±$\pm 2\theta$ are eigenphases of the operator defined in equation (6). Hence, the problem can be solved by applying quantum phase estimation (QPE) to this operator.
@@ II.2 Quantum algorithm
eigenvalue is $\text{eigenvalue is }e^{i2\pi\phi}\rightarrow\text{QPE outputs }r=\phi*K,$
@@
In the case of QAE, we have the eigenphases $\phi=\pm\theta/\pi$. The subspace spanned by $\left\lvert\psi_{0}\right\rangle$ and $\left\lvert\psi_{1}\right\rangle$ finds an orthonormal basis in the two corresponding eigenvectors of the Grover operator. The initial state for QPE can be $\left\lvert psi\right\rangle$. The measured QPE outcome $r$ will be given by:
@@
with probability $8/\pi^{2}$ for $s=1$ and $1-1/(2(s-1))$ for $s$ any other positive integer.
@@ II.3 Classical algorithm
II.4 Alternative quantum algorithms
The quantum algorithm described in section II.2 offers an optimal and straightforward solution to amplitude estimation. However, the required quantum circuits are deep and complex, involving ladders of controlled operations and large numbers of qubits. This renders it unfeasible for current quantum devices, which suffer from noise and - relatedly - limited sizes and continuous computation times.
@@ II.4 Alternative quantum algorithms
A few notable algorithms that follow this structure are phase estimation-free amplitude estimation, or maximum likelihood amplitude estimation (MLAE) [18], faster amplitude estimation (FAE) [20], iterative amplitude estimation [21], and robust amplitude estimation (RAE) [29]. Section II.6 provides an overview of these algorithms.
II.5 Quantum enhanced estimation
@@
For the scope of this work, we are interested in expressions relating the estimation error to the resource count, to be called $\epsilon$ and $N_{q}$ respectively. The former is the root mean squared error, which quantifies the uncertainty, and can be estimated by the standard deviation. The latter quantifies a cost; for instance, a number of measurements, probes or queries, or a (cumulative) evolution time.
@@
II.6 Literature overview
As mentioned in section II.4, the first QFT-free algorithm for QAE was proposed in 2019 by Aaranson and Rall [16], who replaced the QFT by a sequential scheme relying on pure amplitude amplification - that is, on Grover-type circuits, where the only variable is the number of Grover iterations. They call the algorithm ”quantum counting, simplified”; more generally, we use the term”amplitude estimation, simplified” (AES).
@@ II.6 Literature overview
In the same year, other authors had set forth their own alternatives, trading off rigor for practicality. In [18], Suzuki and co-authors proposed a maximum likelihood approach to amplitude estimation (MLAE). Their strategy relies on heuristic sequences of simple Grover circuits, and infers the amplitude based on the measurement data extracted from them. Two heuristics are proposed: LIS and EIS; where the $m$ increase linearly and exponentially respectively. The authors prove lower, but not upper, bounds for the estimation error. The lower bound for EIS is the Heisenberg limit, whereas LIS has a worse (but still quantum-enhanced) best case performance. A numerical analysis shows good performance, although neither of the strategies saturates the lower bound.
@@
The main downside of this approach is the trade-off between computational cost and scalability. To make the problem tractable, the authors work under a Gaussian assumption. This is shown not to significantly affect the numerical results for the case tested therein. While such approximations are often agile for a few parameters, as is the case, they don’t scale well for higher dimensions, which arise when considering more complex noise models.
The statistical representation is arguably the most critical implementation detail in Bayesian inference. It is then important to find methods that are efficient, general and scalable. Failing to do so is bound to frustrate the inference process, affecting not only the acuity of optimality, but correctness itself.
Additionally, RAE requires generalizing the reflections of the Grover operator (6) to arbitrary rotation angles, whereas originally they are fixed and equal to $\pi$. This increases the classical optimization cost, since we are now working with $2m$ parameters (with $m$ the number of circuit layers/applications of the Grover operator) rather than $1$ ($m$ itself).

### v5

@@ Abstract
We present BAE, a problem-tailored and noise-aware Bayesian algorithm for quantum amplitude estimation. In a fault tolerant scenario, BAE is capable of saturating the Heisenberg limit; if device noise is present, BAE can dynamically characterize it and self-adapt. We further propose aBAE, an annealed variant of BAE drawing on methods from statistical inference, to enhance robustness. Our proposals are parallelizable in both quantum and classical components, offer tools for fast noise model assessment, and can leverage preexisting information. Additionally, they accommodate experimental limitations and preferred cost trade-offs. We propose a robust benchmark for amplitude estimation algorithms and use it to test BAE against other approaches, demonstrating its competitive performance in both noisy and noiseless scenarios. In both cases, it achieves lower error than any other algorithm as a function of the cost. In the presence of decoherence, it is capable of learning when other algorithms fail.
1 Introduction
Quantum amplitude estimation (QAE) is a routine for estimating the measurement probability associated with a given subspace for some wavefunction [5]. Notable applications are Monte Carlo integration [35] and estimation tasks [29], with applications in finance [61, 19, 44, 34, 49, 19], chemistry [26, 27, 25]) and machine learning [59, 58, 28, 60].
@@ 1 Introduction
On the other hand, the cost of a naive (deterministic) statistical approximation increases exponentially with the number of parameters. More efficient approximations are vital to ensure tractability; importantly, care must be taken lest they jeopardize scalability or correctness. In [54], a noise-aware approach to QAE based on Bayesian inference is proposed under a normality assumption. While this is a cost-effective approach for simple cases, it is not expected to generalize well to more complex ones, namely problems involving more elaborate noise models. Furthermore, this proposal requires generalizing the Grover operator and optimizing over large numbers of parameters, incurring a significant cost overhead.
1.1 Contributions
@@
In this work, we propose BAE, an approximate Bayesian algorithm for quantum amplitude estimation that is capable of handling noise without making restrictive assumptions or incurring prohibitive costs. Problem tailored heuristics are used to cut down optimization costs while preserving the quantum advantage, bringing the optimization cost per iteration down to constant.
By employing scalable and highly parallelizable statistical methods and other approximations, we further lower the classical processing costs without affecting the quantum advantage. The employed quantum circuits are of the simplest type presented in the literature; they do not require controlled or generalized versions of the Grover operator, nor any other qualitative complication as compared to Grover search.
@@
Our method offers quantifications of model merit with little extra cost, can incorporate available information (such as past experimental data), and is flexible with respect to the cost trade-offs involved. For instance, one may trade in part of the quantum advantage for a lower classical cost, or for increased parallelism in the quantum sub-routine.
@@
To benchmark our algorithm, we numerically test it along with those of other QAE algorithms presented in the literature, and propose thorough, problem-agnostic and cost-effective methods to evaluate and analyze their performance. BAE achieves Heisenberg-limited estimation, has the lowest cost offset among all tested algorithms, and is capable of learning in noisy scenarios where other algorithms fail.
@@
Additionally, inspired by a widely used method in statistics, we propose aBAE, a variant of BAE where the quantum circuits are guided based on measures of statistical efficiency not directly related to the error or uncertainty. This further highlights the importance of this component, and offers a robust and systematic way to work with challenging models or data. We call this variant annealed Bayesian amplitude estimation, after the method it is inspired in, annealed likelihood estimation [47].
The code and datasets used for this paper are available at [42].
1.2 Document structure
@@
The rest of the paper is organized as follows. Section 2 presents relevant background, including the original quantum algorithm, the classical counterpart, and alternative hybrid algorithms. BAE is introduced in section 3. Section 4 describes the numerical methods used for benchmarking. The results of these tests are shown in section 5. Finally, section 6 discusses and summarizes the results, as well as directions for future research. Supplementary details about the algorithm, data processing and background are given in the appendices.
2 Background
2.1 Amplitude estimation
@@
This type of wavefunction arises most commonly in the context of Grover search [22], where given a function $f(x):\{0,1\}^{n}\rightarrow\{0,1\}$, we want to find the pre-image of $1$ under $f$, i.e. $x$ s. t. $f(x)=1$. The operator $\mathrm{\hat{A}}$ is known; we can prepare a state of the form (1), where $\left\lvert\psi_{1}\right\rangle$ is the subspace spanned by solution states and $\left\lvert\psi_{0}\right\rangle$ is its orthogonal complement in the Hilbert space. For that reason, $\left\lvert\psi_{1}\right\rangle$ is often called the good subspace, and $\left\lvert\psi_{0}\right\rangle$ is called the bad subspace. If no prior knowledge is available, the initialization could be a Hadamard transform, assigning equal probability to all possible states. The probability of finding a solution upon measuring state (1) is then $a=2^{-n}$.
@@
2.2 Quantum algorithm
Amplitude estimation is the task of learning $\theta$ or $a$. It can be shown that $\pm 2\theta$ are eigenphases of the operator defined in equation (6). Hence, the problem can be solved by applying quantum phase estimation (QPE) to this operator.
@@ 2.2 Quantum algorithm
$\text{eigenvalue is }e^{i2\pi\phi}\rightarrow\text{QPE outputs }r=\phi*K,$
@@
In the case of QAE, we have the eigenphases $\phi=\pm\theta/\pi$. The subspace spanned by $\left\lvert\psi_{0}\right\rangle$ and $\left\lvert\psi_{1}\right\rangle$ finds an orthonormal basis in the two corresponding eigenvectors of the Grover operator. The initial state for QPE can be $\left\lvert\psi\right\rangle$. The measured QPE outcome $r$ will be given by:
@@
for any positive integer of choice $s$, with probability $8/\pi^{2}$ if $s=1$ and 11/(2(s1))1-1/(2(s-1)) otherwise.
@@ 2.3 Classical algorithm
2.4 Alternative quantum algorithms
The quantum algorithm described in section 2.2 offers an optimal and straightforward solution to amplitude estimation. However, the required quantum circuits are deep and complex, involving ladders of controlled operations and large numbers of qubits. This renders it unfeasible for current quantum devices, which suffer from noise and - relatedly - limited sizes and continuous computation times.
@@ 2.4 Alternative quantum algorithms
A few notable algorithms that follow this structure are phase estimation-free amplitude estimation, or maximum likelihood amplitude estimation (MLAE) [50], faster amplitude estimation (FAE) [36], iterative amplitude estimation [21], and robust amplitude estimation (RAE) [54]. Section 2.6 provides an overview of these algorithms.
2.5 Quantum enhanced estimation
@@
For the scope of this work, we are interested in expressions relating the estimation error to the resource count, to be called $\epsilon$ and $N_{q}$, respectively. The former is the root mean squared error, which quantifies the uncertainty, and can be estimated by the standard deviation. The latter quantifies a cost; for instance, a number of measurements, probes or queries, or a (cumulative) evolution time.
@@
2.6 Literature overview
Key idea Paralle- lizable Circuits Strength(s) Weakness(es) Complexity
QAE [5] QPE on Grover operator no QAA + QPE provably optimal complexity circuit depth and width; noise oblivious
$N_{q}\in\mathcal{O}(\frac{1}{\epsilon}\log\frac{1}{\alpha}\cdot a^{-1})$
MLAE [50] heuristic circuits $\rightarrow$ statistical estimation fully QAA simplicity; numerical performance limited formal guarantees; depth grows indefinitely best case $\mid$ observed: $N_{q}^{\text{(LIS)}}\sim\epsilon^{-0.75}\mid\epsilon^{-0.76}$ $N_{q}^{\text{(EIS)}}\sim\epsilon^{-1}\mid\epsilon^{-0.88}$
V-MLAE [39] MLAE with variational approximation partially QAA numerical per- formance; limits circuit depth limited formal guarantees; cost overhead similar to MLAE except for EIS- observed (untested)
AES [1] rough estimate $\rightarrow$ exponential refining partially QAA provably optimal complexity large quantum cost offset; noise oblivious
$N_{q}\in\mathcal{O}(\frac{1}{\epsilon}\log\frac{1}{\alpha}\cdot a^{-1}$
SAE [56] amplification $\rightarrow$ probability inversion partially Hadamard tests simple non-essential circuits; incon- clusive analysis $N_{q}\in\mathcal{O}(a^{-1})$
IAE [21] watchful optimization of Fisher inf. partially QAA provably near optimal complexity depth grows indefinitely; noise oblivious
M-IAE [15] IAE but distribute shots more favorably partially QAA provably optimal complexity same as IAE
$N_{q}\in\mathcal{O}(\frac{1}{\epsilon}\log\frac{1}{\alpha})$
FAE [36] complementary measurements $\rightarrow$ invert probability partially QAA provably near optimal complexity depth grows indefinitely; noise oblivious
BAE (this paper) problem tailored Bayesian inference partially QAA optimal observed complexity; noise resilient limited formal guarantees; classical cost $N_{q}\in\mathcal{O}(\epsilon^{-1})$ (observed)
As mentioned in section 2.4, the first QFT-free algorithm for QAE was proposed in 2019 by Aaranson and Rall [1], who replaced the QFT by a sequential scheme relying on pure amplitude amplification - that is, on Grover-type circuits, where the only variable is the number of Grover iterations. They call the algorithm ”quantum counting, simplified”; more generally, we use the term”amplitude estimation, simplified” (AES).
@@ 2.6 Literature overview
In the same year, other authors had set forth their own alternatives, trading off rigor for practicality. In [50], Suzuki and co-authors proposed a maximum likelihood approach to amplitude estimation (MLAE). Their strategy relies on heuristic sequences of simple Grover circuits, and infers the amplitude based on the measurement data extracted from them.
Two heuristics are proposed: LIS and EIS; where the $m$ increase linearly and exponentially respectively. The authors prove lower, but not upper, bounds for the estimation error. The lower bound for EIS is the Heisenberg limit, whereas LIS has a worse (but still quantum-enhanced) best case performance. A numerical analysis shows good performance, although neither of the strategies saturates the lower bound. Some authors have studied the behavior of this algorithm in the presence of noise, and suggested improvements [6, 51, 7, 52].
@@
The main downside of this approach is the trade-off between computational cost and scalability. To make the problem tractable, the authors work under a Gaussian assumption. This is shown not to significantly affect the numerical results for the case tested therein. While such approximations are often agile for a few parameters, as is the case, they don’t scale well for higher dimensions, which arise when considering more complex noise models. The statistical representation is arguably the most critical implementation detail in Bayesian inference. It is then important to find methods that are efficient, general and scalable. Failing to do so is bound to frustrate the inference process, affecting not only the acuity of optimality, but correctness itself.
Additionally, RAE requires generalizing the reflections of the Grover operator (6) into arbitrary rotations, whereas originally they are fixed and equal to $\pi$. This increases the classical optimization cost, since we are now working with $2m$ parameters (with $m$ the number of circuit layers/applications of the Grover operator) rather than 11 ($m$ itself).